A construction worker drives to his work 43 km away at an average speed of 62 km/h. Before he gets out of his car he realizes he forgot his hammer at home and drives back at 79 km/h. Assuming his time parked at work was insignificant, what was his average speed overall in km/h?
1
Expert's answer
2020-11-20T09:30:43-0500

By defintion, the average speed is:


"v = \dfrac{total\space distance}{total\space time}"

The total distance is:


"total \space distance =43km+43km = 86km"

The time spent for the way to work is:


"\dfrac{43km}{62km/h} = \dfrac{43}{62}h"

The time spent for the way to home is:


"\dfrac{43km}{79km/h} = \dfrac{43}{79}h"

Thus, the total time is:


"total\space time = \dfrac{43}{62}h + \dfrac{43}{79}h = \dfrac{6063}{4898}h"


And the average speed is:

"v = \dfrac{86km}{\dfrac{6063}{4898}h} \approx 69.5\space km/h"

Answer. "69.5\space km/h".